OpenStudy (anonymous):

what is the solution to the rational equation 3/x-3 = 2/x-2

OpenStudy (anonymous):

\[\frac{3}{x-3}=\frac{2}{x-2}\] use cross multiplication \[3(x-2)=2(x-3)\] \[3x-6=2x-6\] from this it should be fairly straight forward. you can easily check it also

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

x=0